The length of a rectangle 3 feet more than three times its width. The diagonal of the rectangle is 4 feet more than 3 times its width. find the length and width.
Answer by jorel1380(3719)

(3x+4)2-(3x+3)2=x2
9x2+24x+16-9x2-18x-9=x2
0=x2-6x-7
0=(x-7)(x+1)
x=7 or -1
Throwing out the negative answer, we get the width to be 7'. The length is 24'..